OpenStudy (anonymous):

the quadratic equation 3x^2+px+q=0 has roots 2 and -3. find a) value of p and q b) the range of the values of k such that 3x^2+px+q=-3k has no real roots

OpenStudy (anonymous):

sum of the roots=-p/3 product of the roots=q/3 in order to have complex or imaginary roots...discriminant<0.....where discriminant=p^2-4*(3)(q+3k)

OpenStudy (aravindg):

use this : sum of roots =-b/a product of roots =c/a

OpenStudy (anonymous):

then? huh... why i'm so stupid at this?

OpenStudy (maheshmeghwal9):

http://www.learningincontext.com/Algebra-Web/chapt11.htm See for help^_^